OpenStudy (anonymous):

Integrate f(x,y,z)=z over the domain 0<=z<=x^2+y^2<=9

OpenStudy (anonymous):

Domain is \[0 \le z \le x^2 + y^2 \le 9 \]

OpenStudy (anonymous):

I am also convert everything to cylindrical coordinates before I integrate. I am unsure as to what the limits of integration should be. I am thinking that z goes from 0 to r^2 and r goes form 0 to 3

OpenStudy (kainui):

I suggest drawing out the region of integration, that'll help a lot I think in making the other limits of integration perfectly clear.

OpenStudy (kainui):

Another helpful tip, cut up that weird thing of limits of \(0 \le z \le x^2+y^2 \le 9\) into separate pieces: \[0 \le z\]\[z \le x^2+y^2\]\[x^2+y^2 \le 9\]

OpenStudy (anonymous):

Is the region a paraboloid from z=0 to z=9?

OpenStudy (anonymous):

The region is indeed a paraboloid. Suppose we set \(y=0\), then \(z=x^2\) (a parabola in the \(x\)-\(z\) plane); suppose \(x=0\), then \(z=y^2\) (the same but in the \(y\)-\(z\) plane). For any value of \(0\le z\le9\), you get the equation of a circle with radius \(\sqrt z\). See the sketch (not drawn to scale):